# Flat-band FFLO State from Quantum Geometric Discrepancy

### v2

@@ Abstract
Conclusion and Discussion.—In this work, we have identified the QGD as the quantum-geometric origin of the FFLO pairing phase for flat-band electrons. This mechanism differs from other pair-breaking mechanisms, such as the Zeeman effects in helical superconductors Daido et al. (2022) or the orbital effects in layered superconductors Liu (2017). Although we only consider the spin-singlet pairing channel, there is no obstacle in generalizing it to other pseudo-spin degrees of freedom (such as valley and layer) and unconventional pairing channels.
Besides, throughout this work we expect the LO-type pairing state to be more stable than the FF type due to the ±$\pm\bm{Q}$ degeneracy (see Fig. 3) for the inversion-symmetric model Hamiltonian [Eq. (S23)]. By turning on an extra inversion-breaking term, the ±$\pm\bm{Q}$ degeneracy is lifted, resulting in the FF-type pairing state and emergent superconducting diode effect (for more details, see Supplemental Material IVsup ), driven by QGD. This is an analog to the helical superconductors with Rashba spin-orbit interaction and in-plane magnetic fields Smidman et al. (2017); Kinnunen et al. (2018).
Several possible routes can manifest QGD in solid-state systems. For example, optical cavities have been proposed to engineer the topology and quantum geometry of the electronic states Hübener et al. (2021); Schlawin et al. (2022). Another option is that in ultracold atom systems, the hopping parameters can be controlled artificially Jotzu et al. (2015); Kinnunen et al. (2018). These schemes remain the topics for future study.
@@
Acknowledgements—We thank Patrick A. Lee, Xiao-Gang Wen, Xilin Feng, Ying-Ming Xie, Akito Daido, and Qing-Dong Jiang for inspiring discussions. K. T. L. acknowledges the support of the Ministry of Science and Technology, China, and Hong Kong Research Grant Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16307622, and No. 16309223.
@@
Supplemental Material for “Flat-band FFLO State from Quantum Geometric Discrepancy”
Zi-Ting Sun,1 Ruo-Peng Yu,1 Shuai A. Chen,1 Jin-Xin Hu,1 K. T. Law1
1Department of Physics, Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong, China
@@
To study the physics in a multi-band system with pairing instability, it is convenient to project the attractive interaction onto the target flat band. In this section, we will derive the effective action for the single-band description and we will show the full numerical results in the next section. To begin with, we consider a general local attractive Hubbard interaction:
@@
Here $U$ denotes the attractive interaction strength which can give rise to Cooper pairing states. $i$ and $\alpha$ are the site and orbital index, respectively. The fermion operator $c_{\alpha,\bm{k}\sigma}^{\dagger}$ in the momentum space can be expressed by the Fourier transform:
$\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N_{c}}}\sum_{i}e^{-i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{i\alpha\sigma}^{\dagger}$ (S2)
$\hat{c}_{i\alpha\sigma}^{\dagger}$ $=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}.$ (S3)
In this way, we can rewrite the interacting Hamiltonian in the momentum space as
@@
We use $g=U/(N_{orb}N_{c})$ to characterize the effective interaction. We project the interaction onto the target flat band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{a}_{\bm{k}\sigma}$ denotes the band-basis electron annihilation operator. Then from the effective $\hat{H}_{\mathrm{int}}$, on the target flat band, the effective action after the band projection can be expressed as
$S=\sum_{k,\sigma}\bar{a}_{k\sigma}\left(-i\omega_{n}+\xi_{\bm{k}\sigma}\right)a_{k\sigma}-g\sum_{q}\bar{\theta}_{q}\theta_{q}.$ (S6)
where
$\theta_{q}=\sum_{k}\Lambda^{*}(\bm{k},\bm{q})a_{-k\downarrow}a_{k+q\uparrow},$ (S7)
in which $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)=\langle u_{\bm{k}+\bm{q}\uparrow}|u^{*}_{-\bm{k}\downarrow}\rangle$ is introduced as the form factor. Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ being the fermionic (bosonic) Matsubara frequencies.
Figure S1: (a)-(c): The $\Delta-T$ relation for the $\eta=0.1$ (in BCS region) states paired with the finite momenta of $\bm{q}=q(\hat{x}+\hat{y})$, and $q=0$, $0.19\pi$, $0.24\pi$ respectively. The favoured finite-momentum $Q=0$. (d)-(f): The $\Delta-T$ relation for the $\eta=0.23$ (in FFLO region) states paired with the same $q$ in (a)-(c), and the favoured finite-momentum $Q=0.24\pi$. In (a)-(f), $\Delta_{A}$ is donated by the red solid curve and $\Delta_{B}$ is donated by the black dashed curve.
@@
$S_{G}\left[\Delta_{q}\right]=\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (S11)
@@
is the pairing propagator, and the pair susceptibility $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$ at the static limit ($\Omega_{m}=0$) reads
@@
$\chi^{c}_{\bm{q}}=\frac{1}{N_{c}}\sum_{\bm{k}}\frac{1-n_{\mathrm{F}}\left(\xi_{\bm{k+q}\uparrow}\right)-n_{\mathrm{F}}\left(\xi_{-\bm{k}\downarrow}\right)}{\xi_{\bm{k+q}\uparrow}+\xi_{-\bm{k}\downarrow}}|\Lambda(\bm{k},\bm{q})|^{2}.$ (S13)
Near the critical temperature, we also have the number equations (at the weak coupling limit), which read
@@
$N_{\sigma}=-\beta^{-1}\frac{\partial S[\Delta_{q}=0]}{\partial\mu_{\sigma}}=\frac{1}{2}\sum_{\bm{k}}\left[1-\tanh\left(\frac{\beta\xi_{\bm{k}\sigma}}{2}\right)\right].$ (S14)
In the ideal flat-band limit, we have $\xi_{\bm{k}\sigma}=-\mu$. One can find that $\chi^{c}_{\bm{q}}=\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}(1/2-\nu)/[N_{c}\ln(\nu^{-1}-1)]$, where $\nu$ is the filling factor. At half filling $\nu=1/2$, $\chi^{c}_{\bm{q}}\rightarrow\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}/(4N_{c})$.
Figure S2: The superconducting phase diagram of the $\zeta$-lattice model. The red solid line labels $T_{c0}(\eta)$, the black dashed line labels $T_{c\bm{Q}}(\eta)$, where $\bm{Q}=Q(\hat{x}+\hat{y})$. The BCS regime and the FFLO regime are labeled in light blue and yellow, respectively. Parameters used here: $U=0.4t$, $\zeta_{\downarrow}=3$, $W/U=0.1$ in (a) and $W/U=0.2$ in (b).
II II. SELF-CONSISTENT MEAN-FIELD THEORY
@@
In this section, we provide the details of the mean-field numerical calculations in the main text. To be specific, we do the decoupling of the on-site attractive Hubbard interaction within the mean-field approximation
@@
where we have used $\Delta_{\alpha}=-U\left\langle\hat{c}_{\alpha,-\bm{k}\downarrow}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}\right\rangle$ and $\Delta_{\alpha}=\Delta_{\alpha}(\bm{q})$ can be determined by the self-consistent gap equation for a given $\bm{q}$. The full mean-field Hamiltonian can be rewritten in the matrix form in Nambu basis
^$\hat{\mathcal{H}}_{\mathrm{MF}}=\sum_{\bm{k},\bm{q}}\psi^{\dagger}(\bm{k},\bm{q})H_{\mathrm{BdG}}(\bm{k},\bm{q})\psi(\bm{k},\bm{q})+\sum_{\bm{k},\bm{q}}\operatorname{Tr}[H_{-\bm{k}\downarrow}-\mu(\bm{q})\mathbb{I}_{2\times 2}]+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$ (S16)
@@
where $\operatorname{diag}(\Delta_{\alpha})=\operatorname{diag}(\Delta_{A},\Delta_{B})$, the momentum $\bm{q}=q_{x}\hat{x}+q_{y}\hat{y}$, and the Nambu spinor reads
$\psi(\bm{k},\bm{q})=\left(\hat{c}_{A,\bm{k}+\bm{q}\uparrow},\hat{c}_{B,\bm{k}+\bm{q}\uparrow},\hat{c}_{A,-\bm{k}\downarrow}^{\dagger},\hat{c}_{B,-\bm{k}\downarrow}^{\dagger}\right)^{T}.$ (S18)
The grand potential is defined as
@@
In our numerical calculations, we consider the trivial irreducible representation as the intra-orbital pairing channel, i.e., $\Delta_{\alpha\beta}\propto\delta_{{\alpha\beta}}$ is picked up as the only relevant pairing channel, with the other pairing channels neglected. We note that in the previous section, the band-projected interaction Hamiltonian is approximated under the uniform pairing approximation. To show this approximation is accurate near the critical temperature in this model, we plot the $\Delta-T$ relation at different QGD $\eta$ and finite-momentum $\bm{q}$ in Fig. S1. For different $\eta$ and $q$, at low temperatures, $\Delta_{A}$ can differ from $\Delta_{B}$. But near the critical temperature, we find $\Delta_{A}=\Delta_{B}$, which satisfies the uniform pairing condition. This fact explains why our critical temperature analysis based on the uniform pairing condition matches the numerical calculations well.
We also provide more numerical results for the case of finite band dispersion. We keep $W\ll U\ll 2t$ to ensure a narrow bandwidth compared to the band gap and the interaction. In Fig. S2 We show the phase diagram at $W/U=0.1$ and $W/U=0.2$. Compared to Fig.3(b) in the main text, it is clear that the area of the FFLO phase is shrinking with the increase of $W/U$. Especially in Fig. S2(b), the FFLO states can only be found in a very narrow region. It is straightforward to infer that if we keep increasing $W/U$, the FFLO states may vanish, which supports our claim in the main text.
Figure S4: (a) The superconducting phase diagram of $\zeta$-lattice model with an inversion breaking term $\gamma$. The $\bm{Q}>0$ (weak helical superconducting state) and $\bm{Q}<0$ (strong helical superconducting state) regimes are colored in blue and yellow, respectively. Parameters used here: $U=0.4t$, $\zeta_{\uparrow}=3$, $\gamma=3$. (b) The red curve shows the finite-momentum $\bm{Q}=Q(1,1)$ versus QGD ($\eta$). Meanwhile, the blue curve is the corresponding diode efficiency $\Theta$. When increasing $\eta$ to across the phase boundary $\eta=0.26$, the diode efficiency $\Theta$ is enhanced dramatically, and the sign of $\bm{Q}$ is reversed. (c) and (d) The free energy (red curve) and supercurrent (blue curve) of 2 highlighted points in (a) with the temperature $T=0.1T_{0}$ but different $\eta$ [$\eta=0.24$ in (c) and $0.29$ in (d)]. Both of the free energy and supercurrent are normalized to [-1, 1]. The green point is in the $\bm{Q}>0$ phase, and the red point is in the $\bm{Q}<0$ phase.
III III. A FLAT-BAND MODEL ON TRIANGULAR LATTICE
In the main text, we study a $\zeta$-lattice model on a square lattice, which respects $C_{4}$ symmetry. To further show the generality of our approach, in this section, we show similar results for the triangular lattice that has orbitals ($A$ and $B$) per site. Similarly, we can write down the Bloch Hamiltonian as
@@
where $\bm{R}_{1}=a(1,0),\bm{R}_{1}=a(-1/2,\sqrt{3}/2),\bm{R}_{3}=a(-1/2,-\sqrt{3}/2)$. To demonstrate the general validity of our theory in the main text, we calculate the averaged anomalous quantum distance $\bar{d}^{2}_{\eta}(\bm{q})$ for this model in the absence (presence) of QGD in Fig. S3(c) and (d). As a comparison, we also make the plots for the square lattice model presented in the main text in Fig. S3(a) and (b). In all of these figures, the dark blue region denotes the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$. We can see that in both (a) and (c), when $\eta=0$, the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$ is at $\bm{q}=0$. But for (b) and (d), the global minima are not at $\bm{q}=0$, which implies the FFLO instability, as we explained in the main text. It is worth noting that in Fig. S3(b) the FFLO states are four-fold degenerate, while in Fig. S3(d) the FFLO states are six-fold degenerate, due to the different crystalline symmetries of the Bloch function. Our theory shows universality for various model Hamiltonians.
IV. FLAT-BAND HELICAL PAIRING STATE AND SUPERCONDUCTING DIODE EFFECT
The helical superconductivity can be realized in noncentrosymmetric superconductors with the aid of in-plane magnetic field Daido et al. (2022). To achieve an analog of the helical pairing state in a flat-band system with QGD, we manually turn on an extra inversion breaking term in the $\zeta$-lattice model by modifying (recall the periodic function is $\alpha_{\bm{k},\sigma}=\zeta_{\sigma}p_{\bm{k}}$)

### v3

@@ IX Conclusion and Discussion
In this work, we have identified the QGD as a generic quantum-geometric origin of the FFLO pairing phase for flat-band electrons, which can be quantified by AQD. While our analysis focuses on the spin-singlet pairing channel, it can be generalized to other pseudo-spin degrees of freedom (such as valley and layer), unconventional pairing channels, and charge as well as spin channels.
Although we adopt an artificial model here to illustrate the effects of the QGD, we emphasize that Eq. (5) and (6) provide a general perspective on how the QGD delineates FFLO states. Detailed discussions for additional examples of FFLO states with QGD, e.g., inter-valley pairing between different pseudo-Landau levels, can be found in SI III.C sup .
X Acknowledgements
@@
We thank Patrick A. Lee, Xiao-Gang Wen, Kun Yang, Xilin Feng, Ying-Ming Xie, Akito Daido, and Xiao Yan Xu for inspiring discussions. K. T. L. acknowledges the support of the Ministry of Science and Technology, China, and Hong Kong Research Grant Council through Grants No. 2020YFA0309600, No. RFS2021-6S03, No. C6025-19G, No. AoE/P-701/20, No. 16310520, No. 16307622, and No. 16309223.
@@
Supplemental information for “Flat-band FFLO State from Quantum Geometric Discrepancy”
Zi-Ting Sun,1 Ruo-Peng Yu,1 Shuai A. Chen,2 Jin-Xin Hu,3 K. T. Law1
1Department of Physics, Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong, China 2Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Straße 38, Dresden 01187, Germany 3Division of Physics and Applied Physics, School of Physical and Mathematical Sciences, Nanyang Technological University, Singapore 637371
@@
To study the core physics of the superconducting state in a multi-band system, it is convenient to project the attractive interaction onto the target flat band. In this section, we derive the effective action for the single-band description, with the full numerical results presented in the next section.
I.1 A. Linearized gap and number equations
To begin with, we consider a general local attractive Hubbard interaction:
@@
Here $U$ denotes the attractive interaction strength driving the superconducting order. In real space, $i$ and $\alpha$ are the site and orbital index, respectively. The fermion operator $\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}$ in the momentum space can be expressed by Fourier transform via
$\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}=\frac{1}{\sqrt{N_{c}}}\sum_{i}e^{-i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{i\alpha\sigma}^{\dagger},\quad\hat{c}_{i\alpha\sigma}^{\dagger}=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i\alpha}}\hat{c}_{\alpha,\bm{k}\sigma}^{\dagger}.$ (S2)
Therefore, we can rewrite the interacting Hamiltonian in the momentum space as
@@
Here we have used $g=U/(N_{orb}N_{c})$ to characterize the effective interaction. We then project the interaction term onto the target flat band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{a}_{\bm{k}\sigma}$ denotes the band-basis electron annihilation operator. Then from $\hat{H}_{\mathrm{int}}$, the effective action $S$ after the band projection can be expressed as
$S=\sum_{k,\sigma}\bar{a}_{k\sigma}\left(-i\omega_{n}+\xi_{\bm{k}\sigma}\right)a_{k\sigma}-g\sum_{q}\bar{\theta}_{q}\theta_{q},$ (S5)
where $\theta_{q}=\sum_{k}\Lambda^{*}(\bm{k},\bm{q})a_{-k\downarrow}a_{k+q\uparrow}$. Here $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)=\langle u_{\bm{k}+\bm{q}\uparrow}|u^{*}_{-\bm{k}\downarrow}\rangle$ is introduced as the form factor. Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ the fermionic (bosonic) Matsubara frequencies.
@@
$S_{G}\left[\Delta_{q}\right]=S[\Delta_{q}=0]+\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (S9)
@@
is the pairing propagator, and the pairing susceptibility reads
@@
$\chi^{c}_{q}=\frac{1}{N_{c}}\sum_{\bm{k}}\frac{1-n_{\mathrm{F}}\left(\xi_{\bm{k+q}\uparrow}\right)-n_{\mathrm{F}}\left(\xi_{-\bm{k}\downarrow}\right)}{\xi_{\bm{k+q}\uparrow}+\xi_{-\bm{k}\downarrow}-i\Omega_{m}}|\Lambda(\bm{k},\bm{q})|^{2}.$ (S11)
At the static limit ($\Omega_{m}=0$), $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$, and the linearized gap equation is given by $g\chi^{c}_{\bm{q}}=1$.
Near the critical temperature, following the Nosières and Schmitt-Rink (NSR) method De Melo et al. (1993), we consider the Gaussian fluctuations contribution to the thermodynamic potential as $\Omega=\Omega_{0}-\beta^{-1}\sum_{q}\ln\Gamma_{q}$, where $\Omega_{0}=S[\Delta_{q}=0]/\beta$, and the number equation says $N_{\sigma}=-\partial\Omega/\partial\mu_{\sigma}$. We can rewrite $\Omega$ in terms of a phase shift defined by $\Gamma_{q=(\bm{q},i\Omega_{m}\rightarrow\omega\pm i0)}=|\Gamma_{q=(\bm{q},i\Omega_{m}\rightarrow\omega)}|\exp[\pm i\Psi(\bm{q},\omega)]$. So the number equation incorporating the effects of Gaussian fluctuations is given by $N_{\sigma}=N_{0\sigma}+N_{G\sigma}$, where
$N_{0\sigma}=\frac{1}{2}\sum_{\bm{k}}\left[1-\tanh\left(\frac{\beta\xi_{\bm{k}\sigma}}{2}\right)\right],$ (S12)
$N_{G\sigma}=\sum_{\bm{q}}\int_{-\infty}^{+\infty}\frac{d\omega}{\pi}n_{\mathrm{B}}(\omega)\frac{\partial\Psi(\bm{q},\omega)}{\partial\mu_{\sigma}}.$ (S13)
We can introduce the filling factor $\nu_{\sigma}=N_{\sigma}/N_{c}$. In the ideal flat-band limit, we have $\xi_{\bm{k}\sigma}=-\mu_{\sigma}$. When $\nu_{\uparrow}=\nu_{\downarrow}=\nu$ and $\mu_{\uparrow}=\mu_{\downarrow}=\mu$ (no spin-population imbalance), we have $\chi^{c}_{\bm{q}}=\tanh\left(\frac{\beta\mu}{2}\right)\frac{1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})}{2\mu}$. Then we consider the number equation, the free part is $2\nu_{0}=N_{0}/N_{c}=1+\tanh\left(\frac{\beta\mu}{2}\right)$. If we only consider the contribution from $N_{0}$, i.e., neglecting the fluctuations, we can recover Eq. (6) in the main text regardless of the filling factor. However, if we include $N_{G}$, it is not the case for general filling. But worth mention that, Eq. (6) can approximately stand even with fluctuations when $\beta\mu\ll 1$, where $\chi^{c}_{\bm{q}}\rightarrow\beta\left({1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})}\right)/4$. When neglecting $N_{G}$, this condition corresponds to the neighborhood of half filling.
I.2 B. Corrections from Gaussian fluctuations
Then we seriously consider the corrections brought by $N_{G}$, the dynamical pairing propagator at the ideal flat-band limit is
@@
$\Gamma_{q}^{-1}=g^{-1}-\tanh\left(\frac{\beta\mu}{2}\right)\frac{1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})}{2\mu+i\Omega_{m}}.$ (S14)
For the convenience of calculation, we notice that
[…]
If we replace $i\Omega_{m}$ by $\omega+ic$, $c$ is an infinitesimal positive number, we can obtain
[…]
Then from Eq. S15 we have
$\frac{\partial\Psi}{\partial\mu}$ $=\frac{-gc\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\beta}{2\cosh^{2}{\left(\frac{\beta\mu}{2}\right)}\left[\left(2\mu+\omega-g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\right)^{2}+c^{2}\right]}$ (S18)
$+\frac{2gc\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\left(4\mu+2\omega-g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\right)}{\left[\left(2\mu+\omega\right)^{2}+c^{2}\right]\left[\left(2\mu+\omega-g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\right)^{2}+c^{2}\right]},$ (S19)
To evaluate $N_{G}$, we use $\delta(x+a)=\lim_{c\rightarrow 0+}\frac{1}{\pi}\frac{c}{(a+x)^{2}+c^{2}}$ to
$\frac{\partial\Psi}{\partial\mu}=\left[2\pi-\frac{g\pi\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\beta}{2\cosh^{2}{\left(\frac{\beta\mu}{2}\right)}}\right]\delta\left(2\mu+\omega-g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\right)-2\pi\delta\left(2\mu+\omega\right),$ (S20)
so that we arrive at
$N_{G}=-2N_{c}n_{\mathrm{B}}(-2\mu)+\sum_{\bm{q}}\left[2-\frac{g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\beta}{2\cosh^{2}{\left(\frac{\beta\mu}{2}\right)}}\right]n_{\mathrm{B}}\left(-2\mu+g\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})\right)\tanh\left(\frac{\beta\mu}{2}\right)\right).$ (S21)
Figure S1: (a) $T_{c\bm{q}}(\eta)$ from Eq. (6) in the main text (dashed orange line) and from Gaussian fluctuations correction (solid blue line) for $\bm{q}=q(\hat{x}+\hat{y}),\,q\in\left[0,\pi\right]$, at $\eta=0$. (b) Same calculations for $\eta=0.25$. In our calculations, we set $\zeta_{\downarrow}=3$ and $\zeta_{\uparrow}=\zeta_{\downarrow}(1+\eta)/(1-\eta)$, doping factor $\delta\nu=0.1$.
Then we want to show that, even in the presence of Gaussian fluctuations, the FFLO state with the highest critical temperature can still be determined solely by $\bar{d}^{2}_{\mathcal{A}}(\bm{Q})$. We dope the half filling with $\delta\nu$, then the chemical potential becomes $\mu$. Besides, we use the linearized gap equation concerning the FFLO state with finite momentum $\bm{Q}$, namely, $\beta g=4x/\left[\left(1-\bar{d}^{2}_{\mathcal{A}}(\bm{Q})\right)\tanh\left(x\right)\right]$ to replace the corresponding $\beta g$ in the number equation, in which we use $x=\beta\mu/2$ for convenience. Then we reduce the number equation to
[…]
We observe that the solution of $x$ is independent of $g$, which implies the critical temperature is always proportional to $g$. When fixing $\delta\nu$, the solution of $x$ is only determined by the value of $\bar{d}^{2}_{\mathcal{A}}(\bm{Q})$, so does the critical temperature.
However, unlike the free case, $\beta_{c\bm{Q}}$ may not be a monotonic function of $\bar{d}^{2}_{\mathcal{A}}(\bm{Q})$ now. We can plot $\beta_{c\bm{Q}}$-$\bar{d}^{2}_{\mathcal{A}}(\bm{Q})$ curve to look for the global minimum point of $\beta_{c\bm{Q}}$ (highest critical temperature), then we can find out the corresponding FFLO instability with $\bar{d}^{2}_{\mathcal{A}}(\bm{Q})$. We calculate the critical temperatures corrected by Gaussian fluctuations in Fig. S1 as a comparison of Fig. 3(b) in the main text. We can see that the emergence of FFLO instability in the presence of Gaussian fluctuations on the whole aligns with the predictions from QGD and AQD.
II II. SELF-CONSISTENT MULTIORBITAL MEAN-FIELD THEORY
@@
In this section, we provide the detailed methodology for the mean-field numerical simulations discussed in the main text. To be specific, we perform the decoupling of the on-site attractive Hubbard interaction within the mean-field approximation:
@@
where we have used $\Delta_{\alpha}=-U\left\langle\hat{c}_{\alpha,-\bm{k}\downarrow}\hat{c}_{\alpha,\bm{k}+\bm{q}\uparrow}\right\rangle$ and $\Delta_{\alpha}\equiv\Delta_{\alpha}(\bm{q})$ can be determined by the self-consistent gap equation for a given Cooper pair momentum $\bm{q}$. The full mean-field Hamiltonian can be rewritten in the Nambu basis $\psi(\bm{k},\bm{q})=\left(\hat{c}_{A,\bm{k}+\bm{q}\uparrow},\hat{c}_{B,\bm{k}+\bm{q}\uparrow},\hat{c}_{A,-\bm{k}\downarrow}^{\dagger},\hat{c}_{B,-\bm{k}\downarrow}^{\dagger}\right)^{T}$ as
$\hat{\mathcal{H}}_{\mathrm{MF}}=\sum_{\bm{k},\bm{q}}\psi^{\dagger}(\bm{k},\bm{q})H_{\mathrm{BdG}}(\bm{k},\bm{q})\psi(\bm{k},\bm{q})+\sum_{\bm{k},\bm{q}}\operatorname{Tr}[H_{-\bm{k}\downarrow}-\mu(\bm{q})\mathbb{I}_{2\times 2}]+\frac{N_{c}}{U}\sum_{\alpha}|\Delta_{\alpha}|^{2},$ (S24)
@@
where $\operatorname{diag}(\Delta_{\alpha})=\operatorname{diag}(\Delta_{A},\Delta_{B})$ and the momentum $\bm{q}=q_{x}\hat{x}+q_{y}\hat{y}$. The grand potential is defined as
@@
We note that in the previous section, the band-projected interaction Hamiltonian is approximated under the uniform pairing condition, where we only consider the trivial irreducible representation as the intra-orbital pairing channel, such that $\Delta_{\alpha\beta}\propto\delta_{{\alpha\beta}}$ is selected as the only relevant pairing channel, while other pairing channels are neglected. To demonstrate that this approximation is valid near the critical temperature in this model, we plot the $\Delta-T$ relation for different values of the QGD $\eta$ and finite momentum $\bm{q}$ in Fig. S2. At low temperatures, $\Delta_{A}$ and $\Delta_{B}$ can differ depending on $\eta$ and $\bm{q}$. However, near the critical temperature, we observe that $\Delta_{A}=\Delta_{B}$, which satisfies the uniform pairing condition. This observation explains why our critical temperature analysis, based on the uniform pairing condition, aligns well with the numerical results.
We also provide additional numerical results for the case of finite band dispersion. To ensure a narrow bandwidth compared to the band gap and the interaction strength, we maintain the condition $W\ll U\ll 2t$. In Fig. S3 we show the phase diagram for $W/U=0.1$ and $W/U=0.2$. Compared to Fig. 4 in the main text, it is evident that the region corresponding to the FFLO phase shrinks as $W/U$ increases. In particular, Fig. S3(b) shows that the FFLO states are confined to a very narrow region. From this trend, it is reasonable to infer that further increasing $W/U$ could lead to the complete disappearance of the FFLO states. This observation supports the claim made in the main text.
[…]
III III. ADDITIONAL MODELS FOR FFLO PHASE WITH QGD
III.1 A. $\zeta$-lattice model on triangular lattice
In the main text, we investigate a $\zeta$-lattice model on a square lattice, which preserves $C_{4}$ symmetry. To further demonstrate the generality of our approach, we extend our analysis in this section to a triangular lattice featuring two orbitals ($A$ and $B)$ per site. Analogously, we can express the Bloch Hamiltonian as follows
@@
where $\bm{R}_{1}=a(1,0),\bm{R}_{1}=a(-1/2,\sqrt{3}/2),\bm{R}_{3}=a(-1/2,-\sqrt{3}/2)$. We calculate the averaged anomalous quantum distance $\bar{d}^{2}_{\eta}(\bm{q})$ for this model in the absence and presence of QGD, as shown in Fig. S4(c) and (d), respectively. For comparison, we also present the corresponding plots for the square lattice model from the main text in Fig. S4(a) and (b). In all panels, the dark blue regions indicate the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$. We observe that in both (a) and (c), when $\eta=0$ the global minimum of $\bar{d}^{2}_{\eta}(\bm{q})$ occurs at $\bm{q}=0$. However, in (b) and (d), the global minima shift away from $\bm{q}$ implying the FFLO instability, as discussed in the main text. Notice that in Fig. S4(b), the FFLO states exhibit four-fold degeneracy, whereas in Fig. S4(d), they display six-fold degeneracy. This difference arises from the distinct crystalline symmetries of the Bloch functions. These results underscore the universality of our theory across various model Hamiltonians.
III.2 B. Flat-band helical pairing state and superconducting diode effect
Helical superconductivity can be realized in noncentrosymmetric superconductors with the aid of an in-plane magnetic field Daido et al. (2022). To emulate an analogous helical pairing state in a flat-band system with QGD, we introduce an additional inversion-breaking term manually into the $\zeta$-lattice model. This is done by modifying (recall that the periodic function is $\alpha_{\bm{k},\sigma}=\zeta_{\sigma}p_{\bm{k}}$)